OpenStudy (anonymous):

Will give medals. If the angles of a triangle with measures in the ratio of 5:6:7, the find the measure of the smallest angle. a. 40 degrees b. 50 degrees c. 60 degrees d. 70 degrees

OpenStudy (pratyush5):

Let angles be 5x, 6x and 7x Now we know that sum of angles in a triangle is 180 degrees. therefore 5x+ 6x + 7x = 180 You can solve it now.

OpenStudy (anonymous):

But I found the answer already. If someone will check it for me, you can have the medal :) For others: Add the numbers 5 + 6 + 7 = 18 18 will be the denominator. 5/18 6/18 7/18 Multiply these fractions by 180 since all triangles = 180. 5/18 x 180 = 50 6/18 x 180 = 60 7/18 x 180 = 70 These are our angle degrees. 50 is the smallest.

OpenStudy (pratyush5):

You are absolutely right

OpenStudy (pratyush5):

Wow some brains

OpenStudy (pratyush5):

Good work

OpenStudy (anonymous):

Thanks for checking @pratyush5!

OpenStudy (pratyush5):

You're welcome :)